Circular No. 6569

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 960720 AND GRB 970111 J. in ’t Zand, J. Heise, P. Hoyng, and R. Jager, Space Research Or-

ganization of the Netherlands (SRON), Utrecht; L. Piro, E. Costa, and M. Feroci, Istituto Astrofisica Spaziale, CNR, Frascati; F. Frontera, D. Dal Fiume, M. Orlandini, G. Pizzichini, and L. Nicastro, Istituto Tecnologie e Studio delle Radiazioni Extraterrestri, CNR, Bologna; and J. Muller, BeppoSAX, Rome, and SRON, Utrecht, write: “The positions of the xray counterparts to GRB 960720 and GRB 970111 as determined with the BeppoSAX Wide Field Cameras (cf. IAUC 6467 and 6533, respectively) have been refined through an improvement of the spatial calibration of the instruments. The calibration concerns small misalignments between both cameras and between the cameras and the satellite coordinate system, the image scales of both WFCs, and the modeling of the point spread function. The calibration was tested with 123 detections of catalogued x-ray sources with accurately known positions, as found in 15 SAX WFC observations with a variety of orientation angles. These sources have a representative range of intensities and off-axis angles. None of the detections has a difference between catalogue and measured source positions > 3′.9, and 97 percent have a difference of < 3′. In a subset of fourteen sources having significance levels that are within a factor of 2 of the levels of GRB 960720 and GRB 970111, one source shows a difference of > 3′ from the catalogue position. For GRB 960720, we find a 3′-radius error circle centered on α = 17^{h}30^{m}37^{s}, δ = +49^{o}05′.8 (equinox 2000.0). This position is 10′.5 from the previously reported position, for which a 10′ error circle radius was reported. Neither the previously reported SAX NFI source (IAUC 6480) nor the optical source (IAUC 6526) lies within the error circle. The new error circle does include the radio-loud quasar 4C 49.29 (z = 1.038). The error circle for GRB 970111 is centered on α = 15^{h}28^{m}15^{s}, δ = +19^{o}36′.3. This center is 4′.2 from the previously reported position, and the circle overlaps the IPN annulus reported by Hurley et al. (IAUC 6545). None of the previously reported SAX NFI sources (Butler et al., IAUC 6539), RASS sources (Voges et al., IAUC 6539), radio source (Frail et al., IAUC 6545), and ROSAT HRI source (IAUC 6567) lies within the 3′-radius error circle.”

COMET 29P/SCHWASSMANN-WACHMANN 1 Visual m_{1} estimates: Feb. 2.19 UT, 11.6 (K. Hornoch, Lelekovice,

Czech Republic, 0.35-m refl.); 8.15, 14.1: (J. D. Shanklin, Cambridge, England, 0.3-m refractor); 16.18, 14.1 (J. Carvajal, Teruel, Spain, 0.44-m refl.).

1997 March 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green